OpenStudy (anonymous):

If F(x)= x+6 and H(x)= 2X-3 , FIND THE FUNCTION G SUCH THAT G0F=H . PLEASE

OpenStudy (anonymous):

g(x+6) = 2x-3

OpenStudy (anonymous):

really can you please explain it more it it is possible?

OpenStudy (anonymous):

f(x) = x +6 and h(x) = 2x-3 gof (x) = h(x) Plug in the values of f(x) and h(x) g(x +6) = 2x -3 The coefficient of x is 2 in the function and constant term is -3 2(x +6) -15 = 2x -3 So g(x) = 2x -15